mom is 28 years older than zack. mom is 4 years younger than dad. their total age is 84. what is mom's age

Answer by mathman123(1)

zach: x
mom: x+28
dad: x+28+4
x+x+28+x+28+4=84
3x+60=84
3x=84-60
3x=24
x=24/3
x=8
zach=8
mom=8+28= 36
dad=8+28+4= 40